# BEND: Benchmarking DNA Language Models on Biologically Meaningful Tasks

### v1

@@ A.8 Extended results
Model SK-N-SH GM23248 A549 HepG2 HUES64 GM23248 HeLa-S3 ENCFF567KCL ENCFF170XYJ ENCFF948WVD ENCFF690FNR ENCFF890GMD ENCFF840XVU ENCFF754RAW Basset 0.93 0.94 0.93 0.90 0.95 0.94 0.93 CNN 0.84 0.84 0.84 0.82 0.93 0.84 0.83 ResNet-LM 0.86 0.87 0.86 0.85 0.94 0.87 0.86 AWD-LSTM 0.80 0.80 0.80 0.78 0.89 0.80 0.79 NT-H 0.87 0.87 0.87 0.85 0.94 0.87 0.87 NT-MS 0.92 0.92 0.92 0.89 0.96 0.92 0.91 NT-1000G (2.5B) 0.88 0.88 0.88 0.86 0.94 0.88 0.87 NT-V2 0.90 0.91 0.90 0.88 0.96 0.91 0.90 DNABERT 0.91 0.91 0.91 0.88 0.96 0.91 0.90 DNABERT-2 0.89 0.89 0.89 0.87 0.96 0.89 0.89 GENA-LM BERT 0.91 0.91 0.91 0.89 0.95 0.91 0.90 GENA-LM BigBird 0.90 0.91 0.90 0.88 0.95 0.91 0.90 HyenaDNA tiny 0.85 0.85 0.85 0.83 0.92 0.85 0.84 HyenaDNA large 0.91 0.91 0.91 0.88 0.94 0.91 0.90
@@
Model Intron (n=61,229) Intergenic (n=23,379) Upstream gene (n=6,433) Downstream gene (n=4,645) Regulatory region (n=4,049) Noncoding transcript exon (n=3,141) 3’ UTR (n=2,049) 5’ UTR (n=470) TF binding site (n=436) Splice region (n=131) Splice polypyrimidine tract (n=105) Missense (n=62) Splice donor region (n=31) Synonymous (n=24) Splice donor (n=15) Splice donor 5th base (n=13) Stop lost (n=2) DeepSEA 0.7 0.69 0.71 0.71 0.71 0.68 0.64 0.72 0.64 0.55 0.62 0.72 0.46 0.54 0.35 0.83 1.00 ResNet-LM 0.55 0.54 0.56 0.55 0.52 0.51 0.54 0.45 0.46 0.45 0.55 0.51 0.69 0.70 0.50 0.50 1.00 AWD-LSTM 0.51 0.51 0.50 0.52 0.49 0.51 0.52 0.52 0.42 0.52 0.57 0.58 0.62 0.65 0.62 0.55 0.00 NT-H 0.55 0.54 0.55 0.54 0.53 0.52 0.49 0.43 0.44 0.52 0.34 0.58 0.36 0.71 0.50 0.67 1.00 NT-MS 0.55 0.53 0.54 0.55 0.55 0.56 0.57 0.49 0.54 0.54 0.52 0.56 0.56 0.65 0.19 0.60 1.00 NT-1000G-2.5B 0.44 0.43 0.44 0.44 0.48 0.46 0.48 0.44 0.47 0.43 0.42 0.46 0.39 0.54 0.27 0.21 1.00 NT-1000G-500M 0.49 0.48 0.49 0.47 0.50 0.53 0.50 0.46 0.51 0.52 0.50 0.42 0.66 0.29 0.46 0.33 0.00 NT-V2-500M 0.49 0.47 0.47 0.48 0.51 0.48 0.50 0.41 0.51 0.59 0.55 0.36 0.51 0.68 0.77 0.40 0.00 DNABERT 0.60 0.59 0.61 0.60 0.58 0.57 0.55 0.60 0.51 0.52 0.70 0.57 0.77 0.88 0.69 0.45 0.00 DNABERT-2 0.49 0.49 0.48 0.49 0.53 0.48 0.48 0.53 0.58 0.50 0.48 0.57 0.78 0.59 0.35 0.52 1.00 GENA-LM BERT 0.49 0.49 0.50 0.50 0.55 0.51 0.51 0.50 0.55 0.52 0.48 0.55 0.27 0.29 0.58 0.60 0.00 GENA-LM BigBird 0.49 0.48 0.49 0.49 0.53 0.51 0.52 0.50 0.53 0.52 0.48 0.55 0.24 0.43 0.35 0.55 0.00 HyenaDNA large 0.51 0.52 0.50 0.52 0.53 0.51 0.50 0.50 0.48 0.49 0.51 0.49 0.35 0.29 0.19 0.36 0.00 HyenaDNA medium (160k) 0.49 0.50 0.47 0.50 0.52 0.49 0.49 0.48 0.47 0.49 0.51 0.54 0.33 0.31 0.23 0.36 0.00 HyenaDNA medium (450k) 0.50 0.51 0.50 0.51 0.53 0.50 0.50 0.48 0.46 0.48 0.54 0.54 0.39 0.29 0.19 0.43 0.00 HyenaDNA small 0.46 0.47 0.45 0.47 0.50 0.47 0.48 0.47 0.49 0.49 0.48 0.41 0.32 0.25 0.19 0.36 0.00 HyenaDNA tiny 0.47 0.48 0.44 0.49 0.51 0.49 0.48 0.48 0.51 0.49 0.48 0.40 0.34 0.35 0.27 0.24 0.00
@@
Model Intron (n=138,211) Splice region (n=40,360) Splice polypyrimidine tract (n=39,686) Noncoding transcript exon (23,721) 3’ UTR (n=20,441) Splice donor (n=10,811) Splice acceptor (n=9,280) 5’ UTR (n=6,996) Upstream gene (n=2,264) Splice donor region (n=2,056) Splice donor 5th base (n=1,060) Downstream Gene (n=274) Mature miRNA (n=40) Intergenic (n=20) DeepSEA 0.48 0.44 0.46 0.73 0.69 0.47 0.48 0.58 0.61 0.45 0.41 0.72 0.18 0.92 ResNet-LM 0.51 0.67 0.53 0.50 0.51 0.55 0.48 0.46 0.58 0.64 0.63 0.31 0.79 0.05 AWD-LSTM 0.46 0.41 0.47 0.55 0.50 0.51 0.51 0.48 0.47 0.52 0.50 0.44 0.44 0.16 NT-H 0.43 0.53 0.49 0.51 0.56 0.51 0.52 0.53 0.52 0.49 0.50 0.53 0.38 0.58 NT-MS 0.62 0.70 0.65 0.57 0.55 0.74 0.61 0.57 0.56 0.76 0.76 0.44 0.82 0.63 NT-1000G-2.5B 0.49 0.57 0.54 0.52 0.48 0.51 0.50 0.47 0.45 0.52 0.52 0.45 0.10 0.11 NT-1000G-500M 0.46 0.53 0.50 0.49 0.40 0.51 0.49 0.47 0.41 0.47 0.49 0.36 0.03 0.63 NT-V2-500M 0.50 0.52 0.49 0.50 0.36 0.49 0.53 0.52 0.46 0.50 0.43 0.54 0.33 0.21 DNABERT 0.52 0.55 0.47 0.48 0.63 0.54 0.51 0.62 0.58 0.55 0.56 0.62 0.72 0.05 DNABERT-2 0.48 0.46 0.53 0.54 0.49 0.50 0.52 0.45 0.51 0.45 0.52 0.51 0.92 0.95 GENA-LM BERT 0.50 0.49 0.48 0.51 0.50 0.56 0.60 0.50 0.41 0.47 0.42 0.50 0.95 1.00 GENA-LM BigBird 0.48 0.48 0.46 0.51 0.52 0.54 0.60 0.46 0.45 0.48 0.43 0.60 1.00 0.89 HyenaDNA large 0.53 0.52 0.59 0.52 0.48 0.44 0.52 0.48 0.48 0.48 0.42 0.40 0.00 0.63 HyenaDNA medium 160k 0.52 0.54 0.60 0.54 0.46 0.44 0.51 0.46 0.50 0.47 0.41 0.41 0.00 0.47 HyenaDNA medium 450k 0.53 0.53 0.58 0.51 0.46 0.45 0.50 0.51 0.50 0.47 0.42 0.35 0.00 0.37 HyenaDNA small 0.53 0.54 0.59 0.53 0.47 0.44 0.49 0.46 0.53 0.48 0.41 0.44 0.10 0.42 HyenaDNA tiny 0.53 0.55 0.59 0.53 0.47 0.44 0.52 0.48 0.50 0.48 0.42 0.44 0.08 0.37

### v3

@@ A.8 Extended results
Model SK-N-SH GM23248 A549 HepG2 HUES64 GM23248 HeLa-S3 ENCFF567KCL ENCFF170XYJ ENCFF948WVD ENCFF690FNR ENCFF890GMD ENCFF840XVU ENCFF754RAW Basset 0.93 0.94 0.93 0.90 0.95 0.94 0.93 CNN 0.84 0.84 0.84 0.82 0.93 0.84 0.83 ResNet-LM 0.86 0.87 0.86 0.85 0.94 0.87 0.86 AWD-LSTM 0.80 0.80 0.80 0.78 0.89 0.80 0.79 NT-H 0.87 0.87 0.87 0.85 0.94 0.87 0.87 NT-MS 0.92 0.92 0.92 0.89 0.96 0.92 0.91 NT-1000G (2.5B) 0.88 0.88 0.88 0.86 0.94 0.88 0.87 NT-V2 0.90 0.91 0.90 0.88 0.96 0.91 0.90 DNABERT 0.91 0.91 0.91 0.88 0.96 0.91 0.90 DNABERT-2 0.89 0.89 0.89 0.87 0.96 0.89 0.89 GENA-LM BERT 0.91 0.91 0.91 0.89 0.95 0.91 0.90 GENA-LM BigBird 0.90 0.91 0.90 0.88 0.95 0.91 0.90 HyenaDNA tiny 0.85 0.85 0.85 0.83 0.92 0.85 0.84 HyenaDNA large 0.91 0.91 0.91 0.88 0.94 0.91 0.90 GROVER 0.88 0.89 0.88 0.86 0.94 0.89 0.88
@@
Model Intron (n=60,072) Intergenic (n=23,218) Upstream gene (n=6,339) Downstream gene (n=4,581) Regulatory region (n=4,010) Noncoding transcript exon (n=3,099) 3’ UTR (n=2,025) 5’ UTR (n=462) TF binding site (n=433) Splice region (n=129) Splice polypyrimidine tract (n=103) Missense (n=60) Splice donor region (n=31) Synonymous (n=24) Splice donor (n=15) Splice donor 5th base (n=13) Stop lost (n=2) DeepSEA 0.70 0.69 0.71 0.71 0.71 0.68 0.64 0.72 0.64 0.55 0.62 0.72 0.46 0.54 0.35 0.83 1.00 ResNet-LM 0.55 0.54 0.56 0.55 0.52 0.51 0.54 0.44 0.46 0.44 0.54 0.49 0.69 0.70 0.50 0.50 1.00 AWD-LSTM 0.53 0.54 0.52 0.55 0.56 0.53 0.51 0.51 0.51 0.48 0.51 0.44 0.31 0.28 0.38 0.37 0.00 NT-H 0.55 0.54 0.54 0.55 0.52 0.51 0.49 0.43 0.44 0.51 0.33 0.57 0.36 0.71 0.50 0.67 1.00 NT-MS 0.55 0.53 0.54 0.55 0.54 0.55 0.57 0.48 0.53 0.54 0.51 0.54 0.56 0.65 0.19 0.60 1.00 NT-1000G-2.5B 0.44 0.43 0.43 0.44 0.48 0.46 0.48 0.44 0.47 0.42 0.40 0.44 0.39 0.54 0.27 0.21 1.00 NT-1000G-500M 0.49 0.48 0.49 0.47 0.50 0.53 0.50 0.45 0.51 0.51 0.48 0.40 0.66 0.29 0.46 0.33 0.00 NT-V2-500M 0.48 0.47 0.46 0.48 0.50 0.48 0.50 0.41 0.51 0.58 0.54 0.34 0.51 0.68 0.77 0.40 0.00 DNABERT 0.60 0.59 0.61 0.60 0.57 0.57 0.55 0.60 0.51 0.51 0.70 0.57 0.79 0.81 0.65 0.5 0.00 DNABERT-2 0.49 0.49 0.47 0.49 0.53 0.48 0.48 0.52 0.57 0.49 0.47 0.55 0.78 0.59 0.35 0.52 1.00 GENA-LM BERT 0.49 0.49 0.50 0.50 0.54 0.51 0.51 0.49 0.55 0.51 0.47 0.53 0.27 0.29 0.58 0.60 0.00 GENA-LM BigBird 0.49 0.48 0.48 0.49 0.52 0.51 0.52 0.49 0.53 0.51 0.47 0.53 0.24 0.43 0.35 0.55 0.00 HyenaDNA large 0.51 0.52 0.50 0.52 0.53 0.51 0.50 0.49 0.48 0.47 0.54 0.49 0.37 0.26 0.19 0.33 0.00 HyenaDNA medium (160k) 0.48 0.49 0.47 0.50 0.52 0.49 0.49 0.46 0.46 0.49 0.53 0.51 0.34 0.23 0.19 0.33 0.00 HyenaDNA medium (450k) 0.50 0.51 0.49 0.52 0.54 0.50 0.50 0.47 0.45 0.48 0.53 0.54 0.39 0.26 0.19 0.38 0.00 HyenaDNA small 0.46 0.47 0.45 0.47 0.50 0.47 0.48 0.47 0.50 0.49 0.50 0.42 0.32 0.25 0.19 0.33 0.00 HyenaDNA tiny 0.47 0.48 0.44 0.49 0.51 0.49 0.48 0.45 0.50 0.48 0.49 0.39 0.37 0.35 0.23 0.24 0.00 GROVER 0.55 0.55 0.58 0.55 0.55 0.56 0.56 0.50 0.56 0.41 0.48 0.66 0.36 0.46 0.42 0.74 0.00
@@
Model Intron (n=138,211) Splice region (n=40,360) Splice polypyrimidine tract (n=39,686) Noncoding transcript exon (23,721) 3’ UTR (n=20,441) Splice donor (n=10,811) Splice acceptor (n=9,280) 5’ UTR (n=6,996) Upstream gene (n=2,264) Splice donor region (n=2,056) Splice donor 5th base (n=1,060) Downstream Gene (n=274) Mature miRNA (n=40) Intergenic (n=20) DeepSEA 0.48 0.44 0.46 0.73 0.69 0.47 0.48 0.58 0.61 0.45 0.41 0.72 0.18 0.92 ResNet-LM 0.51 0.67 0.53 0.50 0.51 0.55 0.48 0.46 0.58 0.64 0.63 0.31 0.79 0.05 AWD-LSTM 0.53 0.56 0.59 0.52 0.50 0.48 0.52 0.46 0.47 0.50 0.43 0.40 0.12 0.53 NT-H 0.43 0.53 0.49 0.51 0.56 0.51 0.52 0.53 0.52 0.49 0.50 0.53 0.38 0.58 NT-MS 0.62 0.70 0.65 0.57 0.55 0.74 0.61 0.57 0.56 0.76 0.76 0.44 0.82 0.63 NT-1000G-2.5B 0.49 0.57 0.54 0.52 0.48 0.51 0.50 0.47 0.45 0.52 0.52 0.45 0.10 0.11 NT-1000G-500M 0.46 0.53 0.50 0.49 0.40 0.51 0.49 0.47 0.41 0.47 0.49 0.36 0.03 0.63 NT-V2-500M 0.50 0.52 0.49 0.50 0.36 0.49 0.53 0.52 0.46 0.50 0.43 0.54 0.33 0.21 DNABERT 0.52 0.55 0.47 0.48 0.63 0.54 0.51 0.62 0.58 0.55 0.56 0.62 0.72 0.05 DNABERT-2 0.48 0.46 0.53 0.54 0.49 0.50 0.52 0.45 0.51 0.45 0.52 0.51 0.92 0.95 GENA-LM BERT 0.50 0.49 0.48 0.51 0.50 0.56 0.60 0.50 0.41 0.47 0.42 0.50 0.95 1.00 GENA-LM BigBird 0.48 0.48 0.46 0.51 0.52 0.54 0.60 0.46 0.45 0.48 0.43 0.60 1.00 0.89 HyenaDNA large 0.53 0.52 0.59 0.52 0.48 0.44 0.52 0.48 0.48 0.48 0.42 0.40 0.00 0.63 HyenaDNA medium 160k 0.52 0.54 0.60 0.54 0.46 0.44 0.51 0.46 0.50 0.47 0.41 0.41 0.00 0.47 HyenaDNA medium 450k 0.53 0.53 0.58 0.51 0.46 0.45 0.50 0.51 0.50 0.47 0.42 0.35 0.00 0.37 HyenaDNA small 0.53 0.54 0.59 0.53 0.47 0.44 0.49 0.46 0.53 0.48 0.41 0.44 0.10 0.42 HyenaDNA tiny 0.53 0.55 0.59 0.53 0.47 0.44 0.52 0.48 0.50 0.48 0.42 0.44 0.08 0.37 GROVER 0.50 0.46 0.42 0.48 0.54 0.49 0.53 0.55 0.52 0.49 0,45 0.42 0.21 0.21